# Linear stability analysis for a system of singular amplitude equations arising in biomorphology

### v1

@@ 1. Introduction
This system was derived in [WZ3] by multi-scale expansion as a set of amplitude equations formally governing weakly unstable behavior near Turing bifurcation of a general family of convection reaction diffusion system
@@
$u\in\mathbb{R}^{n}$, indexed by bifurcation parameter $\nu$ centered about 00, in the presence of conservation laws: that is, for $R$ of co-rank $0<m\leq n$. As described in [WZ1, WZ3], the main motivation for this study was from modern biomorphology models incorporating mechanical/hydrodynamical effects, in particular vasculogenesis models as in [AGS, SBP]; see also an earlier derivation [HSZ] in the context of Bénard-Marangoni and thin-film flow.
@@
of (1.2), where $\xi=k_{*}(x-d_{*}t)$, $\hat{x}=\varepsilon(x-(d_{*}+\delta)t)$, and $\hat{t}=\varepsilon^{2}t$ for $d_{*},\delta$ real numbers determined by formal expansion. Here, $\nu\sim\varepsilon^{2}$, while $r$ corresponds to the bifurcating neutral direction for the Fourier symbol of the linearized equations about the constant state $u(x,t)\equiv U_{0}$ from which Turing bifurcation occurs. For relations of model parameters of (mcGL) to the form of (1.2), see [WZ3].
@@
denoted the Darcy approximation by analogy to related approximations in hydrodynamics. See, e.g., [JMR] for similar approximations in weakly nonlinear geometric optics and MHD. We note for reference the relations (cf. (1.5))
@@ 1.2. Main results
Main result 1. Our first main result is, under mild genericity assumptions stated in Proposition 4.4 below, the derivation of $m+1$ simple necessary and sufficient conditions for diffusive spectral stability (2.5) of periodic solutions (1.1) of the amplitude equations (mcGL) for $0<\varepsilon\leq\varepsilon_{0}$ sufficiently small, where $m=\dim B$ is the number of mean modes. Here, $\varepsilon_{0}$ may be chosen uniformly on compact parameter-sets for which the above-mentioned genericity conditions are satisfied.
@@
The fact that the order of limits does not affect the value of $\mu_{t}$ can be understood, more generally, from the fact proved in the course of the analysis of Section 4 (regions (ii)-(iii)), that the slow mode $\mu_{t}$ is in fact jointly analytic in $\varepsilon$, $\hat{\sigma}$ for $\varepsilon$, $|\hat{\sigma}|$ sufficiently small, similarly as in the classical case; by contrast, the fast modes $\mu_{c,j}$, as noted in [WZ3] are typically analytic in $\hat{\sigma}$ only on the reduced range $|\hat{\sigma}|\ll\varepsilon$.
@@ 1.3. Discussion and open problems
System (mcGL) generalizes amplitude equations derived for particular examples by Matthews-Cox [MC] in the O(2)-symmetric reaction diffusion case and Häcker-Schneider-Zimmermann [HSZ] in the SO(2)-symmetric reaction convection diffusion case. A major difference in the O(2) case is that the $\varepsilon^{-1}(fB_{\hat{x}}+h|A|^{2}_{\hat{x}})$ convection term does not appear in (mcGL)(ii), cancelling due to reflective symmetry. This leaves a nonsingular reaction diffusion system in place of the coupled reaction diffusion/singular convection diffusion system of (mcGL), for which the analytic issues are essentially different. See [MC, S, WZ3] for further discussion.
@@
Many biomorphology models used in applications feature partial parabolicity and mixed hyperbolic-parabolic type. An important further open problem to extend our results for complete parabolicity to this more general domain, as been done for related settings in [JZN] and references therein. A second very important followup problem is to actually carry out the above-described program for specific biomorphology models, systematically determining associated Turing bifurcations and their stability; that is, studying the practical initiation problem for periodic pattern formation with conservation laws. A third, more speculative followup, as suggested in [WZ1, WZ3], is to study modulation of these patterns as a possible model for emergent dynamics; see [JNRZ, MetZ].
@@ 1.3.1. Relations to thin-film flow
where $\tilde{M}_{1}(\varepsilon)=\tilde{M}_{1}^{0}+\varepsilon\tilde{M}_{1}^{1}$, $\tilde{M}_{1}^{j}$ constant, that is, a singularly perturbed relaxation system with vanishing viscosity $\varepsilon^{2}\partial_{\check{x}}^{2}$, the form ultimately studied in the two-parameter matrix bifurcation analysis of Section 4 It is interesting that a similar relaxation structure was encountered in the studies [JNRZ2, BJNRZ] of a quite different bifurcation occurring in the small-amplitude limit for periodic waves in inclined shallow-water flow in the ultra-small frequency regime, and a reminiscent two-parameter bifurcation analysis successfully carried out. See Remark 4.5 for further discussion.
@@ 2.2.4. Balancing transformation
We notice that this expression vanishes to leading order. In order to match the $O(\varepsilon^{-3})$ term from (2.50)’, we consider the third matrix in (2.49), which has an $O(\varepsilon^{-1})$ coefficient of $\mu_{c}$ coming from the third determinant in (2.49), which gives us
@@
The key quantity $f+2A_{0}hp$ in (2.40) may be regarded as an effective convection obtained by Chapman-Enskog type reduction; note that our equations have the form of a relaxation system.. The condition that $f+2A_{0}hp$ be nonzero is equivalent to noncharacteristicity of this effective convection. Indeed, by the desingularizing rescaling $\delta\to\tilde{\delta}$, we are effectively converting the singular truncated model to a singularly perturbed model
@@ 3. Necessity of Darcy conditions
We next establish necessity of the Darcy stability conditions $\Re(\hat{\sigma})<0$, $\Re\mu_{t}^{0}<0$ determining stability of periodic solutions with respect to the reduced Darcy system (1.18) for stability with respect to the full system (mcGL).
@@ 3.3. Refinement in the scalar case
Hence the Darcy reduction gives a complex Ginzburg-Landau equation of the form studied in [WZ2], from which we can obtain an expression for $\mu_{D,t}$ by either Kato-style expansion as in [WZ2] and then rearranging or by matched asymptotics as done in Appendix B. In either case, the expression we want is
@@ 6.2. Stability
for “tilde model”, a related, higher-order expansion of (mcGL) (denoted the “truncated model” in [WZ2]). Accounting for the difference in truncated vs. tilde models gives an additional $O(\varepsilon)$ error contribution, resulting in (6.7). See [Wh, Appendix A.1], for a detailed computation of errors for the truncated model in an example case with $m=1$. ∎
@@ Appendix A Turing bifurcation for example model
For the model probem (1.4). homogeneous equilibrium states take the form $(\rho_{0},u_{0},c_{0})$, with
@@
On the other hand, for $\alpha\beta=0$, the linearized matrix becomes block lower triangular, and we can solve the dispersion relation explicitly as
@@
We may conclude therefore that a Turing bifurcation occurs for some value $(\alpha\beta)_{*}>0$. Essentially the same argument yields the result for arbitrary $\mu,\nu>0$ as well.
Remark A.1.
It is an interesting question to determine the value of $\alpha\beta$ where bifurcation occurs. In principle this should be determinable by seeking imaginary roots $\lambda=i\tau$ and showing their existence or nonexistence by taking resultants of real and imaginary parts. From the argument above, this occurs strictly before the low-frequency criterion (A.5) is first violated.
Appendix B Matched asymptotics for complex Ginzburg-Landau

### v2

@@ 1. Introduction
The first example of such a system, and the striking phenomenon of singular convection in $B$, was derived via multi-scale expansion by Häcker, Schneider, and Zimmerman [HSZ] in the context of weakly unstable Bénard-Marangoni and thin-film flow, for which $d=0$ and the equations partially decouple. The singular convection they observed is a surprising consequence of the interaction of conserved quantities and convective forces in the underlying pde. For, in the absence of conserved quantities, the relevant amplitude equations are the famous (nonsingular) complex Ginzburg-Landau equations [E, vH, KSM, M3] consisting of the $A$ equation alone, whereas, in the absence of convection, the $B$ equation becomes purely diffusive, and (mcGL) reduces to a nonsingular system discovered previously by Matthews and Cox [MC].
The more general, fully coupled version described above was derived in [WZ3] by multi-scale expansion as a set of amplitude equations formally governing weakly unstable behavior near Turing bifurcation of a general family of convection reaction diffusion system
@@
$u\in\mathbb{R}^{n}$, indexed by bifurcation parameter $\nu$ centered about 00, in the presence of conservation laws: that is, for $R$ of co-rank $0<m\leq n$. As described in [WZ1, WZ3], the main motivation for this study was from modern biomorphology models incorporating mechanical/hydrodynamical effects, in particular vasculogenesis models as in [AGS, SBP].
@@
of (1.2), where $\xi=k_{*}(x-d_{*}t)$, $\hat{x}=\varepsilon(x-(d_{*}+\delta)t)$, and $\hat{t}=\varepsilon^{2}t$ for $d_{*},\delta$ real numbers determined by formal expansion, $c.c.$ denotes complex conjugate, and $H.O.T$ denotes higher-order terms. Here, $\nu\sim\varepsilon^{2}$, while $r$ corresponds to the bifurcating neutral direction for the Fourier symbol of the linearized equations about the constant state $u(x,t)\equiv U_{0}$ from which Turing bifurcation occurs. For relations of model parameters of (mcGL) to the form of (1.2), see [WZ3].
@@
denoted the Darcy approximation by analogy to related approximations in hydrodynamics. See, e.g., [DJMR] for similar approximations in weakly nonlinear geometric optics and MHD. We note for later reference the relations (cf. (1.5))
@@ 1.2. Main results
Main result 1. Our first main result is, under mild genericity assumptions, the derivation of $m+1$ simple necessary and sufficient conditions for diffusive spectral stability (2.5) of periodic solutions (1.1) of the amplitude equations (mcGL) for $0<\varepsilon\leq\varepsilon_{0}$ sufficiently small, where $m=\dim B$ is the number of mean modes. See Proposition 4.4 in the scalar case $m=1$ and Proposition 5.4 in the vector case $m>1$. Here, $\varepsilon_{0}$ may be chosen uniformly on compact parameter-sets for which the above-mentioned genericity conditions are satisfied.
@@
The fact that the order of limits does not affect the value of $\mu_{t}$ can be understood, more generally, from the fact proved in the course of the analysis of Section 4 (regions (ii)-(iii)), that the slow mode $\mu_{t}$ is in fact jointly analytic in $\varepsilon$, $\hat{\sigma}$ for $\varepsilon$, $|\hat{\sigma}|$ sufficiently small, similarly as in the classical case; see Remark 6.5. By contrast, the fast modes $\mu_{c,j}$, as noted in [WZ3] are typically analytic in $\hat{\sigma}$ only on the reduced range $|\hat{\sigma}|\ll\varepsilon$.
@@ 1.3. Discussion and open problems
System (mcGL) generalizes amplitude equations derived for particular examples by Matthews-Cox [MC] in the O(2)-symmetric reaction diffusion case and Häcker-Schneider-Zimmermann [HSZ] in the SO(2)-symmetric reaction convection diffusion case. A major difference in the O(2) case is that the $\varepsilon^{-1}(fB_{\hat{x}}+h|A|^{2}_{\hat{x}})$ convection term does not appear in (mcGL)(ii), canceling due to reflective symmetry. This leaves a nonsingular reaction diffusion system in place of the coupled reaction diffusion/singular convection diffusion system of (mcGL), for which the analytic issues are essentially different. See [MC, S, WZ3] for further discussion.
@@
Many biomorphology models used in applications feature partial parabolicity and mixed hyperbolic parabolic type. An important further open problem to extend our results for complete parabolicity to this more general domain, as been done for related settings in [JZN] and references therein. A second very important followup problem is to actually carry out the above-described program for specific biomorphology models, systematically determining associated Turing bifurcations and their stability; that is, studying the practical initiation problem for periodic pattern formation with conservation laws. In particular, it would be very interesting to carry out a full analysis for the example model discussed in Appendix A. A third, more speculative followup, as suggested in [WZ1, WZ3], is to study modulation of these patterns as a possible model for emergent dynamics; see [JNRZ, MetZ].
@@ 1.3.1. Relations to thin-film flow
where $\tilde{M}_{1}(\varepsilon)=\tilde{M}_{1}^{0}+\varepsilon\tilde{M}_{1}^{1}$, $\tilde{M}_{1}^{j}$ constant, that is, a singularly perturbed relaxation system with vanishing viscosity $\varepsilon^{2}\partial_{\check{x}}^{2}$, the form ultimately studied in the two-parameter matrix bifurcation analysis of Section 4. It is interesting that a similar relaxation structure was encountered in the studies [JNRZ2, BJNRZ] of a quite different bifurcation occurring in the small-amplitude limit for periodic waves in inclined shallow-water flow in the ultra-small frequency regime, and a reminiscent two-parameter bifurcation analysis successfully carried out. See Remark 4.5 for further discussion.
@@ 2.2.4. Balancing transformation
We notice that this expression vanishes to leading order. In order to match the $O(\varepsilon^{-3})$ term from (2.50), we consider the third matrix in (2.49), which has an $O(\varepsilon^{-1})$ coefficient of $\mu_{c}$ coming from the third determinant in (2.49), which gives us
@@
The key quantity $f+2A_{0}hp$ in (2.40) may be regarded as an effective convection obtained by Chapman-Enskog type reduction; note that our equations have the form of a relaxation system. The condition that $f+2A_{0}hp$ be nonzero is equivalent to noncharacteristicity of this effective convection. Indeed, by the desingularizing rescaling $\delta\to\tilde{\delta}$, we are effectively converting the singular truncated model to a singularly perturbed model
@@ 3. Necessity of Darcy conditions
We next establish necessity of the Darcy stability conditions $\Re(\hat{\sigma})<0$, $\Re\mu_{t}^{0}<0$, determining stability of periodic solutions with respect to the reduced Darcy system (1.18), for stability with respect to the full system (mcGL).
@@ 3.3. Refinement in the scalar case
Hence the Darcy reduction gives a complex Ginzburg-Landau equation of the form studied in [WZ2], from which we can obtain an expression for $\mu_{D,t}$ by either Kato-style expansion as in [WZ2] and then rearranging or by matched asymptotics as done in Appendix C. In either case, the expression we want is
@@ 6.2. Stability
for the “tilde model”, a related, higher-order expansion of (mcGL) (denoted the “truncated model” in [WZ2]). Accounting for the difference in truncated vs. tilde models gives an additional $O(\varepsilon)$ error contribution, resulting in (6.7). See [Wh, Appendix A.1], for a detailed computation of errors for the truncated model in an example case with $m=1$. ∎
@@ Appendix A Turing bifurcation for example model
For the model problem (1.4). homogeneous equilibrium states take the form $(\rho_{0},u_{0},c_{0})$, with
@@
On the other hand, for $\alpha\beta=0$, the linearized matrix becomes block triangular, and we can solve the dispersion relation explicitly as
@@
We may conclude therefore that a Turing bifurcation occurs for some value $(\alpha\beta)_{*}>0$. Essentially the same argument yields the result for arbitrary $\mu,\nu>0$ as well. It would be very interesting to determine the nature and properties of this bifurcation, in the spirit of the current analysis.
Appendix B Numerical illustration, case $m=1$
In the below figures, we display the results of a numerical comparison of the spectra of the linearized equations for (mcGL) and the results predicted by Taylor expansion, for a generic example system with $m=1$. The parameters chosen are $a=1+i$, $b=1$, $c=-3+2i$, $d=-1+2i$, $e_{B}=1$, $f=1$, $g=2+2i$, $h=2$, $\varepsilon=10^{-}2$. In the illustrations, blue dots denote the real parts of (numerically approximated) true eigenvalues, green, predictions from $\mu_{t}$, and red, predictions from $\mu_{c}$, with lefthand panel displaying results for small $\sigma$ and righthand large $\sigma$, and each pair of panels associated with a different wave-number $\kappa$.
In the left column, we’ve plotted the curves on $\lvert\hat{\sigma}\rvert\leq 10\varepsilon$ and in the right column, we have the curves on $\lvert\hat{\sigma}\rvert\leq 1$. We note the strong agreement between $\mu_{c}\hat{\sigma}^{2}$ and $\Re\lambda_{c}(\hat{\sigma})$ on $\hat{\sigma}=o(\varepsilon)$ and good agreement between $\mu_{t}\hat{\sigma}^{2}$ and $\Re\lambda_{t}(\hat{\sigma})$ on the region where $\hat{\sigma}=o(1)$, as is expected by Remark 6.5. We emphasize that the singularity in the dispersion relations are quite prominent, as $\Re\lambda_{c}(\hat{\sigma})$ reaches $O(1)$ when $\hat{\sigma}\sim\varepsilon$, in contrast to the classical Ginzburg-Landau or Matthews-Cox cases where $\Re\lambda_{c}(\hat{\sigma})\sim\varepsilon^{2}$ when $\hat{\sigma}\sim\varepsilon$. Note also that the effects of singularity are visible in the right-hand column of figures through the “spikes” near $\hat{\sigma}=0$. The final point to which we wish to draw emphasis is that it is $\lambda_{c}$ and $\lambda_{s}$ that intersect and lose analyticity within an $O(\varepsilon)$ domain, while $\lambda_{t}$ remains analytic on a domain of $o(1)$. We recall that this played a key role in the analysis of Cases (ii)-(iv). One final remark is that, for fixed $\varepsilon$, the agreement between $\mu_{t}\hat{\sigma}^{2}$ and $\lambda_{t}(\hat{\sigma})$ gets worse as $\lvert\kappa\rvert$ increases. This is to be expected, as the formula for $\mu_{t}$ in (2.52) goes to $\infty$ as $\lvert\kappa\rvert^{2}\to\kappa_{E}^{2}$.
[…]
Figure 1. Wavenumber $\kappa=0$.
[…]
Figure 2. Wavenumber $\kappa=\kappa_{E}/4$.
[…]
Figure 3. Wave number $\kappa=\kappa_{E}/2$.
Appendix C Matched asymptotics for complex Ginzburg-Landau